The Ujsag Printing Co. wants to calculate the cost of adding a third machine. Comment on the following statement made by the company accountant: “If we add a third printing press, our total costs will increase from $1060 to $1191. Therefore the marginal cost of the third machine is $1191/3=$397.” Is the accountant correct? Explain your reasoning.
Answer #1

No, the accountant is not correct. This is because what the accountant have calculated, i.e., $1191/3=$397 is the Average Cost of the 3rd unit and not the Marginal Cost. Marginal Cost of third machine = Change in Total Cost by adding third machine = 1191 - 1060 = $131. Thus, the marginal cost of third machine is $131.